OpenStudy (anonymous):

there were 37,000 people at a ball game in LA. the days receipts were $262000. how many people paid $11 for reserved seats and how many paid $6 for general admission?

OpenStudy (saifoo.khan):

R+G=37,000 OR R=37,000-G 11(37,000-G)+6G=262,000 407,000-11G+6G=262,000 -5G=262,000-407,000 -5G=-145,000 G=-145,000/-5 G=29,000 ANS. FOR THE NUMBER OF GENERAL ADMISSION TICKETS SOLD. 37,000-29,000=8,000 RESERVED SEAT TICKETS SOLD.

OpenStudy (anonymous):

thanks dude

OpenStudy (saifoo.khan):

np